OpenStudy (anonymous):

help me solve this? x/4 = x-2/3

Parth (parthkohli):

Try to multiply 4 to both sides.

Parth (parthkohli):

\( \color{Black}{\Rightarrow x = 4(x - {2\over 3})}\)

Parth (parthkohli):

The next step is to distribute 4 amongst x and -2/3

Parth (parthkohli):

\( \color{Black}{\Rightarrow x = 4x - {8 \over 3}}\) \( \color{Black}{\Rightarrow x - 4x = -{8 \over 3}}\) \( \color{Black}{\Rightarrow -3x = -{8 \over 3} }\) \( \color{Black}{\Rightarrow x = \Large {-8 \over 3} \div -3}\) \( \color{Black}{\Rightarrow x = \Large {-8 \over 3} \times {-1 \over 3} }\)

OpenStudy (anonymous):

or multiply both sides by 12 to remove fraction. it would be easier.

Parth (parthkohli):

\( \color{Black}{\Rightarrow x = \Large {8 \over 9}}\)

Parth (parthkohli):

@taylorneedsmathhelp Did you get it? :)

Parth (parthkohli):

Do you need any more help?

OpenStudy (anonymous):

yes

Parth (parthkohli):

Sure, ready to help. What is your question? Can you post it after closing this one?

OpenStudy (anonymous):

i just dont understand how to to get the answer to this, like what do i do with the 8/9 & what do you mean post after closing this one?

Parth (parthkohli):

x = 8/9 This is the value of x.

OpenStudy (anonymous):

okay but my options dont even say 8/9 they say, a. 4 b. 8 c.2

OpenStudy (anonymous):

[if \alpha and \beta are the zeroes of the given polynomialf(x)=5x ^{2}-7x+1,find the value of 1/\alpha and 1/\]

Parth (parthkohli):

@s.aishwarya Can you please delete this and make your own post?

OpenStudy (anonymous):

yaar abhi issi main answer kar do plz later i will definately do plzzz

OpenStudy (anonymous):

did u get the answer